OpenStudy (shamil98):

A hot bowl of soup cools according to Newton's law of cooling, what is the initial temperature? The function where t is time (in minutes ) is T(t) = 68 + 144e^-.04t I just need to confirm my first step. The first step is to set t=0 for the initial temperature right?

OpenStudy (shamil98):

T(0) = 68 + 144e^-0.4(0) T(0) = 68 + 144 T(0) = 212